OpenStudy (anonymous):

two numbers whose product is 64 and the sum is 8

OpenStudy (solomonzelman):

\(\Huge\color{green}{ \rm a \times b =64 }\) \(\Huge\color{green}{ \rm a + b =8 }\) solve, doesn't matter which one is which.

OpenStudy (solomonzelman):

NO

OpenStudy (solomonzelman):

Look at what you wrote again.

OpenStudy (solomonzelman):

So, you see what your mistake is ?

OpenStudy (anonymous):

let 2 numbers be x and y xy=64 sum = f(x) = x+y = x+64/x now minimize f(x)

OpenStudy (anonymous):

Yes,

OpenStudy (solomonzelman):

Anyway, ... \(\Huge\color{red}{ \it a + b =8 }\) \(\Huge\color{red}{ \it a =8-b }\) \(\Huge\color{red}{ \it (b-8) \times b =64 }\) \(\Huge\color{red}{ \it b^{2}-8b =64 }\) \(\Huge\color{red}{ \it b^{2}-8b +64=0 }\) solve using the quadratic formula, ....

OpenStudy (solomonzelman):

the last equation is \(\Huge\color{red}{ \it b^{2}-8b -64 =0 }\)